The ultraviolet rays that cause sunburn is called UV rays. They have a wavelength of


approximately 300 nanometers (3.0 x 10-7 m). What is the frequency of a UV ray?


1
Expert's answer
2021-11-11T17:08:35-0500
"f=\dfrac{c}{\lambda}=\dfrac{3\times10^8\ \dfrac{m}{s}}{3\times10^{-7}\ m}=1.0\times10^{15}\ Hz."